I need to solve this equation by finding square roots. I don't really understand how to do the problem. I think the answer is 8.94427191, but I don't understand why. The problemis:
x^2+8=72
Please help me.

Answer by checkley71(8403)

X^2+8=72
X^2=72-8
X^2=64
X=SQRT64
X=8 ANSWER.
PROOF:
8^2+8=72
64+8=72
72=72